OpenStudy (anonymous):

what is the domain of the function y=4sqrt2x+6

OpenStudy (jdoe0001):

well, just bear in mind that a negative number inside a square root, gives you an imaginary value, so, what they're really asking is What values can "x" take inside the square root that will not make it give a negative number

OpenStudy (jdoe0001):

0 is ok, \(\large \sqrt{0}=0\) so that's ok

OpenStudy (jdoe0001):

so, what do you think?

OpenStudy (anonymous):

\[y=4\sqrt{2x+6}\] a.\[x \ge 3\] b.\[x \ge -3\] c. \[x > \] d.\[x \ge \frac{ -1 }{ 3 }\]

OpenStudy (jdoe0001):

so, what do you think? which values can "x" have without yielding a negative value inside the root?

OpenStudy (anonymous):

i dont know how to do this very well, square roots have alwasy confused me

OpenStudy (jdoe0001):

ok, lemme put it differently, let's say \(\sqrt{2x+6}=\sqrt{0} meaning 2x+6=0 if 2x+6=0, what would be the value of "x"?

OpenStudy (jdoe0001):

woops

OpenStudy (jdoe0001):

ok, lemme put it differently, let's say \(\large \sqrt{2x+6}=\sqrt{0}\) meaning 2x+6=0 if 2x+6=0, what would be the value of "x"?

OpenStudy (anonymous):

well it would have to be a negative value right?

OpenStudy (jdoe0001):

dunno

OpenStudy (anonymous):

-6 right?

OpenStudy (jdoe0001):

let's see, 2x+6=0 .... what is "x" there?

OpenStudy (jdoe0001):

well, 2x=-6, then divide both by 2

OpenStudy (anonymous):

-3?

OpenStudy (jdoe0001):

right, so 2x+6=0 WHEN X=-3, so if "x" goes BELOW that number, -3, 2(-4)=-8, ==> -8+6=-2 or \(\large \sqrt{-2}\), which is a no-dice for a square root

OpenStudy (anonymous):

so b.?

OpenStudy (jdoe0001):

so, for the function to "work", "x" has to be ABOVE that number, -3 and that's your answer

OpenStudy (jdoe0001):

yes, b) :)

OpenStudy (anonymous):

so \[x \ge -3\]

OpenStudy (anonymous):

thank you can i have help with a couple more?

OpenStudy (jdoe0001):

sure, just post them in the channel, so we all can see it and help and revise each other :)

OpenStudy (anonymous):

what is the domain function \[y=4\sqrt{2x-5}\] a. \[x= \ge \frac{ 5 }{ 2 }\] b.\[x \ge \frac{ 2 }{ 5 }\] c.\[x \ge -\frac{ 5 }{ 2 }\] d.\[x> -\frac{ 5 }{ 2 }\]

OpenStudy (jdoe0001):

do the same procedure when they talk about the DOMAIN, they mean what values can the "independent variable" can take and yield a valid expression in the case of roots, for EVEN roots, the number inside it, CANNOT BE negative

OpenStudy (jdoe0001):

so, find out where the expression = 0, and find "x" from there, then you know when "x" is THAT VALUE, the expression is 0, and if "x" goes BELOW OR ABOVE that, it'd be negative. so once you know where the expression is 0, and what "x" is, use the next number for "x" and see what the expression gives

OpenStudy (anonymous):

it would be - 2.5? right?

OpenStudy (jdoe0001):

so, when this expression = 0, what's "x"?

OpenStudy (anonymous):

which would be -5/2 which would be d? right?

OpenStudy (jdoe0001):

well, "x" can be equal to \(\large \frac{5}{2}\) , at that value you get 0 and \(\large \sqrt{0}=0\)

OpenStudy (jdoe0001):

x>SOMETHING, means "x" HAS TO BE greater than SOMETHING

OpenStudy (jdoe0001):

\(\cfrac{-5}{2}\) I meant :|

OpenStudy (anonymous):

so am i wrong? im sorry im realllllyyy bad at math

OpenStudy (jdoe0001):

"x" can be equal to \(\large \cfrac{-5}{2}\) , it can't go below that though, it can go above though

OpenStudy (anonymous):

so its d?

OpenStudy (jdoe0001):

see the differences? :)

OpenStudy (anonymous):

ik that... its a???

OpenStudy (jdoe0001):

yes, it'd be "a" because that says "x" CAN BE EQUAL TO \(\large \cfrac{-5}{2}\) OR GREATER, and the root would still be non-negative

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

what is sin A fot the triangle shown

OpenStudy (jdoe0001):

on that picture, the angle being used is "c" you could say, you're EYEING from "c" for your triangle, put your EYE on "a", who is your opposite and who is your hypotenuse?